OpenStudy (anonymous):

What value of k makes the equation true? 12 + (–17) = –17 + k k =

OpenStudy (anonymous):

.. i dont get it

OpenStudy (anonymous):

@SydtheKid913

OpenStudy (anonymous):

hello?

OpenStudy (sydthekid913):

hi sorry am making a power point on the history of automobiles :P

OpenStudy (sydthekid913):

So what we have here is the associative property of addition.

OpenStudy (sydthekid913):

now you see that on one side we have 12 + (-17) and then we have an = sign now what does = mean?

OpenStudy (sydthekid913):

= means equivalent or the same as so if we have 12 + (-17) is the same as -17 + k then we see that both sides of the equation must be the same

OpenStudy (anonymous):

it means equals the same

OpenStudy (sydthekid913):

so what is 12 + (-17) ?

OpenStudy (anonymous):

-5

OpenStudy (sydthekid913):

that is correct now what do we have to do to -17 to get it to equal -5?

OpenStudy (anonymous):

+12?

OpenStudy (sydthekid913):

correct!!! great job! :D

OpenStudy (anonymous):

yay

OpenStudy (sydthekid913):

do you see why your answer would be 12? it is because the equation was 12 + (-17) = -17 + k so if they are equal then you know you have to get one side to equal the other side You probably already got that but I just want to make sure :P :D

OpenStudy (anonymous):

yea.

OpenStudy (sydthekid913):

lol XD

OpenStudy (sydthekid913):